In a test of reaction time, we want to know if there is a difference in the speed at which people click on small vs. large buttons. 10 participants are presented with large button sizes and reaction times are measured (in milliseconds). Then 10 different participants, matched in age, are presented with small button sizes, then reaction times are measured. Use an alpha of .05. See below table for summary data:

Mean Standard Deviation
Large Buttons 681.60 35.22
Small Buttons 756.80 32.13
Difference ?? 28.78

State the hypotheses.Set rejection criteria. Compute the test statistic (i.e., standardized score). Make a statistical determination.
Answer #1

X, speed at which people click on small button. X, speed at which people click on Targe botton & x = 681.60 ni=10 S₂ = 32. 13= 28.78 0.1to.. = 28.78 s t = soliti Jn 12 / 1 t 10 V 10 = 28.78 (0.4472) o - 12.871 S.E. test statistic t = x - x2 = 681.60P value o Decision Reject Ho if p valve za Here Pvalue = 0 L 2 =0.05 i we reject Ho. Conchesion we may conclede that there is